jhonyy9 (jhonyy9):

- so if we need to prove that every primes grater than 2 can be writing in the form of 2n+1 - how is possible do it this proof ?

OpenStudy (anonymous):

well, proving is quite perspiring job mentally..

jhonyy9 (jhonyy9):

in mathematics this not working acceptably never

OpenStudy (anonymous):

Suppose that p is prime and that p > 2. Then if \[p \neq 2n + 1\] for some n, then p must be even. Hence it can be written in the form p=2n for some n. However p is prime by assumption, and cannot be divided by anything but itself and 1. 2n can be divided by 2 which is a contradiction. Hence p is of the form p=2n+1 for some n. Lol I'm not convinced by this but it's been a long time since i've seen any arguments like this...

OpenStudy (anonymous):

callum29, could you help my problem??

OpenStudy (anonymous):

please??

jhonyy9 (jhonyy9):

,,Callum29" - thank you - so this is right sure and understandably easy but in mathematics every statement is acceptably when this was proven that is true ... - so can you do it , can you prove it ?

OpenStudy (anonymous):

now I don't understand your question...

OpenStudy (anonymous):

can you help me with this callum29?? please? x=2(mod 3) x=3(mod 7) x=7(mod 5) find a unique incongruent solution to modulo n. (number theory problem using chinese remainder theorem) please help!!

jhonyy9 (jhonyy9):

,,Callum29" can you prove it that every primes grater than 2,can be writing in the form of 2n+1 ?